Voice Control Company has some issued shares for which the announced dividend is $7 and the dividend will grow at a rate of 6% for the next 5 years (Year1-5). From year 6 the growth rate will change to 8% and will continue thereafter. What is the current price of the share if the required rate of return rate is 10%?

 


1
Expert's answer
2021-12-14T14:33:17-0500

The divident in Year 5 is:

"D5 = 7\u00d71.06^5 = 9.37."

The current price of the share is:

"P = \frac{9.37} {0.1 - 0.08} = 468.5."